OpenStudy (anonymous):

TRUE OR FALSE .7 + x = x + 7 is an example of the Identity Property of Addition.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

5+6=6+5

OpenStudy (anonymous):

False...commutative

OpenStudy (anonymous):

7--2 doesnt = -2-7

OpenStudy (anonymous):

"commutative" might be right but i know that the statement 7+x=x+7 is a true satement,now i understand the question

OpenStudy (anonymous):

That means its...False, its not the identity. LOL

OpenStudy (anonymous):

yeah u right

OpenStudy (anonymous):

False they are not equal!